Compute the total and annual return on the following investment.

Two years after paying ​$3000 for shares in a startup​ company, you sell the shares for $2400 ​(at a​ loss).

The total return is    %

​(Do not round until the final answer. Then round to the nearest tenth as​ needed.)

The annual return is. %.
Answer #1

Given that

Two years after paying ​$3000 for shares in a startup​ company, you sell the shares for $2400 ​(at a​ loss).

the total return is = (2400-3000)/3000 = -20%

the annual return is = (2400/3000)^(1/2) = 0.8944 = 89.44%

Kindly Upvote if Helpful :)

HOPE THIS MAY HELP YOU----------------------

------------------------------THANK YOU---------------------------